OpenStudy (chlobug):

Hi, everyone! I just finished a warm-up and instruction for my next Chemistry unit. However, I am having trouble with this homework-assignment question. A sample of tin (Cp = 0.227 J/gC) is placed in a freezer. Its temperature decreases from 15.0 to 10.0C as it releases 543 J of energy. What is the mass of the sample? Round your answer to three significant figures. Please help me understand this; thanks!

OpenStudy (aaronq):

\(\sf energy~(heat)=mass*specific~heat~capacity*change~in~temperature\)

OpenStudy (chlobug):

Thanks. I'll use that equation. Would it be: 543 J = m x 25.0 C x 5 C?

OpenStudy (aaronq):

almost, the specific heat capacity has the symbol \(C_p\), look at your question again

OpenStudy (chlobug):

Okay thanks. I'll try it again. Thanks for bearing with me! 543 J = m x 0.227 x 5

OpenStudy (aaronq):

that's right! no problem dont forget to round

OpenStudy (chlobug):

I cannot thank you enough! I appreciate your help. I'll definitely give you a medal! If I have another question in Chemistry, will you give me permission to please tag your name? (:

OpenStudy (chlobug):

Also, I used that equation and this was my work: 543 = m x .227 x 5 543 = m x 1.135 543 + 1.135 = 478.41 m = 478.41 The actual answer was 95.7g. What did I do wrong? Thanks (:

OpenStudy (aaronq):

you messed up the algebra \(543 = m * 1.135\) \(\dfrac{543}{1.135} = m \)

OpenStudy (chlobug):

\[543\div1.135=478.41\] Right? How do I get the 95.7 from there? Thanks (: I'm sorry for all of the questions!

OpenStudy (aaronq):

oh, sorry i thought you just did the algebra wrong. that is the answer

OpenStudy (chlobug):

No problem. Thank you. Unfortunately, the actual answer was 95.7. I'm not sure what I did wrong.

OpenStudy (aaronq):

It's not. you must've made an error copying it down or the book youre getting this off made a mistake. :S

OpenStudy (chlobug):

Okay. I'll ask my teacher tomorrow (: Thank you for your time, @aaronq! You've been awesome.

OpenStudy (aaronq):

no problem! hope you get this cleared up